OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (konradzuse):

How do you tihnk we start this. In your own words ^^^^

OpenStudy (anonymous):

distributive property?

OpenStudy (konradzuse):

You do use the property, though I don't think you need to explain each property, just the steps... Properties are deff a good bonus tho.

OpenStudy (konradzuse):

That is the correct first step I would do though, now showit please.

OpenStudy (anonymous):

so its gonna be 3-5x-10=3x+3+6 -7-5x=3x+9 add 5 both side -7=8x+9 substract 9 both side -16=8x divide 8 both side and the answer is: x=-2

OpenStudy (konradzuse):

There you go :)

OpenStudy (anonymous):

yeee :) thxx